OpenStudy (anonymous):

Write the following equation in slope-intercept form. 3x - 2y = 5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

translate as "solve for \(y\) "

OpenStudy (misty1212):

subtract \(3x\) from both sides, divide by \(-2\)

OpenStudy (anonymous):

thx misty. :)

OpenStudy (hockeychick23):

ok so slope intercept form is y= mx+b so first we need to get y by itself. So: 3x - 2y = 5 -->3x - 2y-3x = 5-3x Now your problem looks like this: -2y= -3x+5 Next you divide both sides by -2 y= 3/2x-5/2

OpenStudy (anonymous):

thx hockeychick. :)

OpenStudy (hockeychick23):

no problem!(: